# Equivalence Relations on Vertex Operator Algebras, II: Witt Equivalence and Orbifolds

### v1

@@ 6.1. Relationship to RCFT Classification
Assuming 2.21 (or more directly 2.22) and that 5.6 holds for strongly rational vertex operator algebras $V^{\prime}$ with $[\operatorname{Rep}(V^{\prime})]=[\operatorname{Vect}]$, it follows that any positive, strongly rational vertex operator algebra $V$ can be obtained by starting with $\mathsf{E}_{8,1}^{\otimes n}\cong V_{E_{8}^{n}}$ for some sufficiently large $n\in\mathbb{Z}_{\geq 0}$, performing a sequence of orbifolds and then taking a commutant.
@@
Moving backwards, it follows from 2.22 that $V$ can be embedded into a holomorphic vertex operator algebra $U$ and obtained via the coset construction as $V=\operatorname{Com}_{U}(W)$, where $W=\operatorname{Com}_{U}(V)$. On the other hand, $U$ and $\mathsf{E}_{8,1}^{\otimes n}$ are Witt equivalent for $n\coloneqq c(U)/8\in\mathbb{Z}_{\geq 0}$ and satisfy $[\operatorname{Rep}(U)]=[\operatorname{Rep}(\mathsf{E}_{8,1}^{\otimes n})]=[\operatorname{Vect}]$; by assumption, they can be connected by a sequence of orbifolds. ∎
@@
The proof above reveals that we do not require the full might of 5.6. We only need to demand that it hold for vertex operator algebras with $[\operatorname{Rep}(V^{\prime})]=[\operatorname{Vect}]$. But we showed in 5.14 that this is true when $c(V^{\prime})\leq 24$, at least if one accepts the uniqueness of the moonshine module, 2.20.
@@
Assuming 2.21 (or more directly 2.22) and that 5.8 holds for RCFTs ${{}_{V^{\prime}}}\mathcal{H}^{\prime}_{W^{\prime}}$ with $\operatorname{Rep}(V^{\prime})$ being Witt trivial, it follows that any positive RCFT ${{}_{V}}\mathcal{H}_{W}$ can be obtained by starting with $\mathsf{E}_{8,1}^{\otimes n}\otimes\overline{\mathsf{E}}_{8,1}^{\otimes m}$ for some large enough $n,m\in\mathbb{Z}_{\geq 0}$, performing a sequence of orbifolds and then quotienting decoupled degrees of freedom.
@@
From 2.22, it follows that $V$ and $W$ can be embedded into holomorphic vertex operator algebras (or chiral CFTs) $A$ and $B$, respectively. Define $V^{\prime}=\operatorname{Com}_{A}(V)$ and $W^{\prime}=\operatorname{Com}_{B}(W)$, which are both strongly rational vertex operator algebras. Since $\operatorname{Rep}(V)$ is Witt equivalent to $\operatorname{Rep}(W)$ and since $\operatorname{Rep}(V^{\prime})\cong\overline{\operatorname{Rep}(V)}$ and $\operatorname{Rep}(W^{\prime})\cong\overline{\operatorname{Rep}(W)}$, it follows that $\operatorname{Rep}(V^{\prime})$ is Witt equivalent to $\operatorname{Rep}(W^{\prime})$. In particular, one can form an RCFT of the form ${{}_{V^{\prime}}}\mathcal{H}^{\prime}_{W^{\prime}}$. The tensor product ${{}_{V}}\mathcal{H}_{W}\otimes{{}_{V^{\prime}}}\mathcal{H}^{\prime}_{W^{\prime}}$ can be thought of as an RCFT with respect to the chiral algebras $V\otimes V^{\prime}$ and $W\otimes W^{\prime}$. That is, there is some $\mathcal{H}^{\prime\prime}$ such that
@@ 6.2. Relationship to Deformation Classes of QFTs
Assume 5.7, and suppose we are given an RCFT ${{}_{V}}\mathcal{H}_{W}$ for which we aim to construct a boundary. We learn from 4.10 that there is a simple topological line interface $I$ of finite quantum dimension between the gapless chiral boundaries $V$ and $W$ on which $\mathcal{H}$ terminates, as on the right of Figure 14. By folding this picture as if closing a book, one obtains a boundary condition $B$ with finite $g$-function of the RCFT ${{}_{V}}\mathcal{H}_{W}$. Thus, 6.3 holds.
Conversely, suppose that any RCFT admits a boundary condition with finite $g$-function, and suppose $V$ and $W$ are two Witt equivalent strongly rational vertex operator algebras. Choose any topological interface $\mathcal{H}$ between $\operatorname{Rep}(V)$ and $\operatorname{Rep}(W)$ and form the RCFT ${{}_{V}}\mathcal{H}_{W}$. By assumption, this RCFT has a boundary condition. One can inflate this to a three-dimensional Kapustin–Saulina picture as on the right of Figure 14 to obtain a topological line interface $I$ between $V$ and $W$ on which $\mathcal{H}$ terminates. This means that $V$ and $W$ are interface equivalent, and hence that 5.7 is true. ∎
@@
Figure 14. A three-dimensional depiction of a boundary condition $B$ of an RCFT ${{}_{V}}\mathcal{H}_{W}$.
@@ 7. Witt Equivalence and Generalized Symmetries
In this section, we discuss the relationship between Witt equivalence and generalized global symmetries. In Section 7.1, we relate the problem of classifying vertex operator algebras within a weak Witt class to the problem of understanding symmetries. In Section 7.2, we refine this idea further by describing how a generalization of quantum Galois theory enters into the picture. In Section 7.3, we describe an approach to classifying generalized global symmetries in RCFTs. We conclude with examples in Section 7.4. In particular, we give a physical argument that all the finite symmetries of the $\mathrm{SU}(2)_{1}$ Wess–Zumino–Witten model are invertible, and we also find a Fibonacci symmetry of the moonshine module.
@@ 7.1. Symmetry-Subalgebra Duality
Indeed, note that $V$ defines a Lagrangian algebra $L(V)$ in $\operatorname{Rep}(W)$, and therefore a topological boundary condition in the TQFT $(\operatorname{Rep}(W),c(W))$. [^15] More pedantically, since the 3d TQFT has non-vanishing central charge, what we call a “topological boundary condition” should be more properly thought of as a topological interface to the invertible TQFT with the same central charge. Consider placing $(\operatorname{Rep}(W),c(W))$ on an interval, with its $W$ boundary imposed on one end and the topological boundary condition $L(V)$ imposed on the other. The topological boundary condition supports a fusion category worth of topological line operators $\operatorname{Ver}(V/W)\coloneqq\operatorname{Rep}(W)_{L(V)}$, which is the category of $L(V)$-modules inside the modular tensor category $\operatorname{Rep}(W)$. On the other hand, by dimensionally reducing along the interval direction, we recover the original theory $V$, and immediately see that $\operatorname{Ver}(V/W)$ can be thought of as furnishing a symmetry of $V$. Moreover, as advertised, this three-dimensional perspective reveals that the local operators in $W$ are transparent to the lines in $\operatorname{Ver}(V/W)$, by virtue of the fact that the latter are supported on the topological boundary while the former live on the $W$ boundary (which is often called the “physical” boundary in the literature on symmetry TQFTs). See Figure 15.
@@ 7.4. Examples
From the three-dimensional perspective afforded by Figure 15, $\mathsf{D}_{8,1}$ has the toric code (i.e. $\mathbb{Z}_{2}$-gauge theory) as its bulk TQFT. Labeling the anyons of the toric code as $\mathbf{1},e,m,f$, the following identification with modules of $\mathsf{D}_{8,1}$ holds,
@@
$(Z(\operatorname{Vect}_{\mathbb{Z}_{2}}),8)=\{\mathsf{D}_{8,1}\},\;(Z(\operatorname{Vect}_{\mathbb{Z}_{2}}^{\omega}),8)=\{\mathsf{A}_{1,1}\mathsf{E}_{7,1}\},\;(Z(\operatorname{Fib}),8)=\{\mathsf{G}_{2,1}\mathsf{F}_{4,1}\}.$
@@
There has been ongoing work to identify a strongly rational, $c=8$ vertex operator algebra $V$ that has the Drinfeld center $Z(\operatorname{Hg})$ of the Haagerup fusion category $\operatorname{Hg}$ as its representation category, i.e. to show that the symbol $(\mathcal{C},c)=(Z(\operatorname{Hg}),8)$ corresponds to a valid bulk genus, and to construct a vertex operator algebra representative of this genus. For example, it is known that if such a theory exists, then its character vector $\smash{\operatorname{ch}_{V(\alpha)}(\tau)=\operatorname{tr}_{V(\alpha)}q^{L_{0}-1/3}}$ must be one of two options, calculated in [EG11]. From 7.2, we learn that the existence of such a vertex operator algebra $V$ would imply that $\mathsf{E}_{8,1}$ has a $\operatorname{Hg}$ symmetry satisfying $\mathsf{E}_{8,1}^{\operatorname{Hg}}\cong V$.
@@
As another example, we use symmetry-subalgebra duality and quantum Galois theory to study a particular non-invertible symmetry of the moonshine module $V^{\natural}$. This chiral CFT is celebrated for its invertible symmetries, which furnish the simple sporadic monster group, $\smash{\operatorname{Aut}(V^{\natural})}\cong\mathbb{M}$. However, the full category of topological line defects contains (infinitely) many non-invertible symmetries as well. See, e.g., [LS21] for an example of an action of the Tambara–Yamagami fusion category $\mathcal{TY}_{+}(\mathbb{Z}_{2})\cong\operatorname{Rep}(L(1/2))$ on $V^{\natural}$.
We investigate actions of the unitary Fibonacci category $\operatorname{Fib}$ on $V^{\natural}$. [^18] While this work was being completed, the paper [FEH24] appeared on the arXiv, which has overlapping content. By symmetry-subalgebra duality, if such a symmetry exists, then the operators of the monster CFT that are transparent to the Fibonacci line will participate in a strongly rational conformal subalgebra $W\coloneqq\smash{(V^{\natural})^{\operatorname{Fib}}}$ with $\operatorname{Rep}(W)=Z(\operatorname{Fib})\cong\operatorname{Fib}\boxtimes\overline{\operatorname{Fib}}$. Conversely, we may prove the existence of a Fibonacci symmetry by finding such a conformal subalgebra $W$. It follows from the results of [Ray24] (see Section E.4.18) that there are only two possibilities for the character vector $\operatorname{Ch}_{W}$ of such a putative $W$ if it is to satisfy positivity (and hence, only two possibilities for the Fibonacci twisted sector partition functions of $V^{\natural}$):
(7) $\operatorname{Ch}_{W}(\tau)=\begin{pmatrix}[l]q^{-1}+59045q+6092960q^{2}+241395800q^{3}+5623284096q^{4}+\dots\\ 8671q^{3/5}+1804672q^{8/5}+97010251q^{13/5}+2713483488q^{18/5}+\dots\\ q^{-3/5}+1568q^{2/5}+672918q^{7/5}+45506688q^{12/5}+\dots\\ 137839q+15400800q^{2}+622904170q^{3}+\dots\\ \end{pmatrix}$
@@
(8) $\operatorname{Ch}_{W}(\tau)=\begin{pmatrix}[l]q^{-1}+55920q+5961710q^{2}+239095800q^{3}+5597184096q^{4}+\dots\\ 2q^{-2/5}+7299q^{3/5}+1731552q^{8/5}+95550116q^{13/5}+\dots\\ 1938q^{2/5}+693294q^{7/5}+45905184q^{12/5}+1437360744q^{17/5}+\dots\\ 140964q+15532050q^{2}+625204170q^{3}+14648672160q^{4}+\dots\\ \end{pmatrix}.$
We shall show that at least the first option is realized in the monster.
In [HLY12], it was shown that every element of the 3A conjugacy class of $\mathbb{M}$ fixes a $\mathbb{Z}_{3}$-parafermion vertex operator algebra $\mathcal{P}_{3}=L(4/5,0)\oplus L(4/5,3)\subset\smash{V^{\natural}}$ (see [MR24, LABEL:MR1ex:para] for the definition). Also, the commutant $\smash{\textsl{VF}_{24}^{\natural}=V^{\natural}/\mathcal{P}_{3}}$ was shown to possess an action of the Fischer sporadic group $\mathrm{Fi}_{24}$. The tensor product $\smash{\mathcal{P}_{3}\otimes\textsl{VF}_{24}^{\natural}}$ occurs as a strongly rational conformal subalgebra of $\smash{V^{\natural}}$ with representation category given by $\smash{\operatorname{Rep}(\mathcal{P}_{3}\otimes\textsl{VF}_{24}^{\natural})\cong Z(\operatorname{Vect}_{\mathbb{Z}_{3}}\boxtimes\operatorname{Fib})}$, and $\smash{V^{\natural}}$ is obtained from $\smash{\mathcal{P}_{3}\otimes\textsl{VF}_{24}^{\natural}}$ by condensing the canonical Lagrangian algebra corresponding to the Dirichlet boundary condition. Thus, the subalgebra $\smash{\mathcal{P}_{3}\otimes\textsl{VF}_{24}^{\natural}\subset V^{\natural}}$ defines a $\smash{\operatorname{Ver}(V^{\natural}/\mathcal{P}_{3}\otimes\textsl{VF}_{24}^{\natural})}=\operatorname{Vect}_{\mathbb{Z}_{3}}\boxtimes\operatorname{Fib}$ symmetry of $\smash{V^{\natural}}$, with the $\operatorname{Vect}_{\mathbb{Z}_{3}}$ living in the 3A conjugacy class of $\mathbb{M}$.
Let us study the predictions of quantum Galois theory. 7.3 says that the vertex operator subalgebras of $\smash{V^{\natural}}$ containing $\smash{\mathcal{P}_{3}\otimes\textsl{VF}_{24}^{\natural}}$ are in bijection with fusion subcategories of $\operatorname{Vect}_{\mathbb{Z}_{3}}\boxtimes\operatorname{Fib}$. Besides $\operatorname{Vect}_{\mathbb{Z}_{3}}\boxtimes\operatorname{Fib}$ and $\operatorname{Vect}$, which correspond to $\smash{\mathcal{P}_{3}\otimes\textsl{VF}_{24}^{\natural}}$ and $\smash{V^{\natural}}$, respectively, there are the fusion subcategories $\operatorname{Vect}_{\mathbb{Z}_{3}}$ and $\operatorname{Fib}$. We call the corresponding vertex operator subalgebras predicted by 7.3 $U$ and $W$, respectively. There are thus the following Hasse diagrams:

### v2

@@ 6.1. Relationship to RCFT Classification
Assuming 2.22 (or more directly 2.23) and that 5.6 holds for strongly rational vertex operator algebras $V^{\prime}$ with $[\operatorname{Rep}(V^{\prime})]=[\operatorname{Vect}]$, it follows that any positive, strongly rational vertex operator algebra $V$ can be obtained by starting with $\mathsf{E}_{8,1}^{\otimes n}\cong V_{E_{8}^{n}}$ for some sufficiently large $n\in\mathbb{Z}_{\geq 0}$, performing a sequence of orbifolds and then taking a commutant.
@@
Moving backwards, it follows from 2.23 that $V$ can be embedded into a holomorphic vertex operator algebra $U$ and obtained via the coset construction as $V=\operatorname{Com}_{U}(W)$, where $W=\operatorname{Com}_{U}(V)$. On the other hand, $U$ and $\mathsf{E}_{8,1}^{\otimes n}$ are Witt equivalent for $n\coloneqq c(U)/8\in\mathbb{Z}_{\geq 0}$ and satisfy $[\operatorname{Rep}(U)]=[\operatorname{Rep}(\mathsf{E}_{8,1}^{\otimes n})]=[\operatorname{Vect}]$; by assumption, they can be connected by a sequence of orbifolds. ∎
@@
The proof above reveals that we do not require the full might of 5.6. We only need to demand that it hold for vertex operator algebras with $[\operatorname{Rep}(V^{\prime})]=[\operatorname{Vect}]$. But we showed in 5.14 that this is true when $c(V^{\prime})\leq 24$, at least if one accepts the uniqueness of the moonshine module, 2.21.
@@
Assuming 2.22 (or more directly 2.23) and that 5.8 holds for RCFTs ${{}_{V^{\prime}}}\mathcal{H}^{\prime}_{W^{\prime}}$ with $\operatorname{Rep}(V^{\prime})$ being Witt trivial, it follows that any positive RCFT ${{}_{V}}\mathcal{H}_{W}$ can be obtained by starting with $\mathsf{E}_{8,1}^{\otimes n}\otimes\overline{\mathsf{E}}_{8,1}^{\otimes m}$ for some large enough $n,m\in\mathbb{Z}_{\geq 0}$, performing a sequence of orbifolds and then quotienting decoupled degrees of freedom.
@@
From 2.23, it follows that $V$ and $W$ can be embedded into holomorphic vertex operator algebras (or chiral CFTs) $A$ and $B$, respectively. Define $V^{\prime}=\operatorname{Com}_{A}(V)$ and $W^{\prime}=\operatorname{Com}_{B}(W)$, which are both strongly rational vertex operator algebras. Since $\operatorname{Rep}(V)$ is Witt equivalent to $\operatorname{Rep}(W)$ and since $\operatorname{Rep}(V^{\prime})\cong\overline{\operatorname{Rep}(V)}$ and $\operatorname{Rep}(W^{\prime})\cong\overline{\operatorname{Rep}(W)}$, it follows that $\operatorname{Rep}(V^{\prime})$ is Witt equivalent to $\operatorname{Rep}(W^{\prime})$. In particular, one can form an RCFT of the form ${{}_{V^{\prime}}}\mathcal{H}^{\prime}_{W^{\prime}}$. The tensor product ${{}_{V}}\mathcal{H}_{W}\otimes{{}_{V^{\prime}}}\mathcal{H}^{\prime}_{W^{\prime}}$ can be thought of as an RCFT with respect to the chiral algebras $V\otimes V^{\prime}$ and $W\otimes W^{\prime}$. That is, there is some $\mathcal{H}^{\prime\prime}$ such that
@@ 6.2. Relationship to Deformation Classes of QFTs
Assume 5.7, and suppose we are given an RCFT ${{}_{V}}\mathcal{H}_{W}$ for which we aim to construct a boundary. We learn from 4.11 that there is a simple topological line interface $I$ of finite quantum dimension between the gapless chiral boundaries $V$ and $W$ on which $\mathcal{H}$ terminates, as on the right of Figure 15. By folding this picture as if closing a book, one obtains a boundary condition $B$ with finite $g$-function of the RCFT ${{}_{V}}\mathcal{H}_{W}$. Thus, 6.3 holds.
Conversely, suppose that any RCFT admits a boundary condition with finite $g$-function, and suppose $V$ and $W$ are two Witt equivalent strongly rational vertex operator algebras. Choose any topological interface $\mathcal{H}$ between $\operatorname{Rep}(V)$ and $\operatorname{Rep}(W)$ and form the RCFT ${{}_{V}}\mathcal{H}_{W}$. By assumption, this RCFT has a boundary condition. One can inflate this to a $3$-dimensional Kapustin–Saulina picture as on the right of Figure 15 to obtain a topological line interface $I$ between $V$ and $W$ on which $\mathcal{H}$ terminates. This means that $V$ and $W$ are interface equivalent, and hence that 5.7 is true. ∎
@@
Figure 15. A $3$-dimensional depiction of a boundary condition $B$ of an RCFT ${{}_{V}}\mathcal{H}_{W}$.
@@ 7. Witt Equivalence and Generalized Symmetries
In this section, we discuss the relationship between Witt equivalence and generalized global symmetries. In Section 7.1, we relate the problem of classifying vertex operator algebras within a weak Witt class to the problem of understanding symmetries. In Section 7.2, we refine this idea further by describing how a generalization of quantum Galois theory enters into the picture. In Section 7.3, we describe an approach to classifying generalized global symmetries in RCFTs. We conclude with examples in Section 7.4. In particular, we give a physical argument that all the finite symmetries of the $\mathrm{SU}(2)_{1}$ Wess–Zumino–Witten model are invertible, and we also find two Fibonacci symmetries of the moonshine module.
@@ 7.1. Symmetry-Subalgebra Duality
Indeed, note that $V$ defines a Lagrangian algebra $L(V)$ in $\operatorname{Rep}(W)$, and therefore a topological boundary condition in the TQFT $(\operatorname{Rep}(W),c(W))$. [^16] More pedantically, since the 3d TQFT has non-vanishing central charge, what we call a “topological boundary condition” should be more properly thought of as a topological interface to the invertible TQFT with the same central charge. Consider placing $(\operatorname{Rep}(W),c(W))$ on an interval, with its $W$ boundary imposed on one end and the topological boundary condition $L(V)$ imposed on the other. The topological boundary condition supports a fusion category worth of topological line operators $\operatorname{Ver}(V/W)\coloneqq\operatorname{Rep}(W)_{L(V)}$, which is the category of $L(V)$-modules inside the modular tensor category $\operatorname{Rep}(W)$. On the other hand, by dimensionally reducing along the interval direction, we recover the original theory $V$, and immediately see that $\operatorname{Ver}(V/W)$ can be thought of as furnishing a symmetry of $V$. Moreover, as advertised, this $3$-dimensional perspective reveals that the local operators in $W$ are transparent to the lines in $\operatorname{Ver}(V/W)$, by virtue of the fact that the latter are supported on the topological boundary while the former live on the $W$ boundary (which is often called the “physical” boundary in the literature on symmetry TQFTs). See Figure 16.
@@
Finally, we note that 7.2 can be generalized to Witt classes $([\mathcal{C}],c)$ with $[\mathcal{C}]$ different from $[\operatorname{Vect}]$ using the results of [Rie22].
@@ 7.4. Examples
From the $3$-dimensional perspective afforded by Figure 16, $\mathsf{D}_{8,1}$ has the toric code (i.e. $\mathbb{Z}_{2}$-gauge theory) as its bulk TQFT. Labeling the anyons of the toric code as $\mathbf{1},e,m,f$, the following identification with modules of $\mathsf{D}_{8,1}$ holds,
@@
$(Z(\operatorname{Vect}_{\mathbb{Z}_{2}}),8)=\{\mathsf{D}_{8,1}\},\;(Z(\operatorname{Vect}_{\mathbb{Z}_{2}}^{\omega}),8)=\{\mathsf{A}_{1,1}\mathsf{E}_{7,1}\},\;(Z(\mathrm{Fib}),8)=\{\mathsf{G}_{2,1}\mathsf{F}_{4,1}\}.$
@@
There has been ongoing work to identify a strongly rational, $c=8$ vertex operator algebra $V$ that has the Drinfeld center $Z(\mathrm{Hg})$ of the Haagerup fusion category $\mathrm{Hg}$ as its representation category, i.e. to show that the symbol $(\mathcal{C},c)=(Z(\mathrm{Hg}),8)$ corresponds to a valid bulk genus, and to construct a vertex operator algebra representative of this genus. For example, it is known that if such a theory exists, then its character vector $\smash{\operatorname{ch}_{V(\alpha)}(\tau)=\operatorname{tr}_{V(\alpha)}q^{L_{0}-1/3}}$ must be one of two options, calculated in [EG11]. From 7.2, we learn that the existence of such a vertex operator algebra $V$ would imply that $\mathsf{E}_{8,1}$ has a $\mathrm{Hg}$ symmetry satisfying $\mathsf{E}_{8,1}^{\mathrm{Hg}}\cong V$.
@@
As another example, we use symmetry-subalgebra duality and quantum Galois theory to study non-invertible symmetries of the moonshine module $V^{\natural}$. This chiral CFT is celebrated for its invertible symmetries, which furnish the simple sporadic monster group, $\smash{\operatorname{Aut}(V^{\natural})}\cong\mathbb{M}$. However, the full category of topological line defects contains (infinitely) many non-invertible symmetries as well. See [LS21, Vol24], for examples of Tambara-Yamagami categories [TY98] acting on $V^{\natural}$.
Here, we investigate actions of the unitary Fibonacci category $\mathrm{Fib}$ on $V^{\natural}$. [^19] While this work was being completed, the paper [FEH24] appeared on the arXiv, which has overlapping content. By symmetry-subalgebra duality, if such a symmetry exists, the operators of the monster CFT that are transparent to the Fibonacci line participate in a strongly rational conformal subalgebra $W\coloneqq\smash{(V^{\natural})^{\mathrm{Fib}}}$ with $\operatorname{Rep}(W)=Z(\mathrm{Fib})\cong(G_{2},1)\boxtimes(F_{4},1)$. Conversely, we may prove the existence of a Fibonacci symmetry by finding such a conformal subalgebra $W$. It follows from the results of [Ray24] (see Section E.4.18) that there are only two possibilities for the character vector $\operatorname{Ch}_{W}$ of such a putative vertex operator algebra $W$ if it is to satisfy positivity (and hence, only two possibilities for the Fibonacci twisted sector partition functions of $V^{\natural}$):
(7) $\operatorname{Ch}_{W}(\tau)=\begin{pmatrix}[l]q^{-1}+59045q+6092960q^{2}+241395800q^{3}+\dots\\ 8671q^{3/5}+1804672q^{8/5}+97010251q^{13/5}+\dots\\ q^{-3/5}+1568q^{2/5}+672918q^{7/5}+45506688q^{12/5}+\dots\\ 137839q+15400800q^{2}+622904170q^{3}+\dots\\ \end{pmatrix}$
@@
(8) $\operatorname{Ch}_{W}(\tau)=\begin{pmatrix}[l]q^{-1}+55920q+5961710q^{2}+239095800q^{3}+\dots\\ 2q^{-2/5}+7299q^{3/5}+1731552q^{8/5}+95550116q^{13/5}+\dots\\ 1938q^{2/5}+693294q^{7/5}+45905184q^{12/5}+\dots\\ 140964q+15532050q^{2}+625204170q^{3}+14648672160q^{4}+\dots\\ \end{pmatrix}.$
We shall show that both options are realized by conformal subalgebras of the moonshine module, and hence that there are at least two Fibonacci topological line defects in $V^{\natural}$ that cannot be conjugated into each other using invertible symmetries. It would be interesting to prove that these two are the only Fibonacci actions on $V^{\natural}$ up to conjugation.
Option 1. First, let us describe how to construct a Fibonacci action of the moonshine module corresponding to (7).
In [HLY12], it was shown that every element of the 3A conjugacy class of $\mathbb{M}$ fixes a $\mathbb{Z}_{3}$-parafermion vertex operator algebra $\mathcal{P}(3)=L(\nicefrac{{4}}{{5}})\oplus L(\nicefrac{{4}}{{5}},3)\subset\smash{V^{\natural}}$ (see [MR24, LABEL:MR1ex:para] for the definition). Also, the commutant $\smash{\textsl{VF}_{24}^{\natural}=V^{\natural}/\mathcal{P}(3)}$ was shown to possess an action of the Fischer sporadic group $\mathrm{Fi}_{24}$. The tensor product $\smash{\mathcal{P}(3)\otimes\textsl{VF}_{24}^{\natural}}$ occurs as a strongly rational conformal subalgebra of $\smash{V^{\natural}}$ with representation category given by $\smash{\operatorname{Rep}(\mathcal{P}(3)\otimes\textsl{VF}_{24}^{\natural})\cong Z(\operatorname{Vect}_{\mathbb{Z}_{3}}\boxtimes\mathrm{Fib})}$, and $\smash{V^{\natural}}$ is obtained from $\smash{\mathcal{P}(3)\otimes\textsl{VF}_{24}^{\natural}}$ by condensing the canonical Lagrangian algebra corresponding to the Dirichlet boundary condition. Thus, the subalgebra $\smash{\mathcal{P}(3)\otimes\textsl{VF}_{24}^{\natural}\subset V^{\natural}}$ defines a $\smash{\operatorname{Ver}(V^{\natural}/\mathcal{P}(3)\otimes\textsl{VF}_{24}^{\natural})}=\operatorname{Vect}_{\mathbb{Z}_{3}}\boxtimes\mathrm{Fib}$ symmetry of $\smash{V^{\natural}}$, with the $\operatorname{Vect}_{\mathbb{Z}_{3}}$ living in the 3A conjugacy class of $\mathbb{M}$.
Let us study the predictions of quantum Galois theory. 7.3 says that the vertex operator subalgebras of $\smash{V^{\natural}}$ containing $\smash{\mathcal{P}(3)\otimes\textsl{VF}_{24}^{\natural}}$ are in bijection with fusion subcategories of $\operatorname{Vect}_{\mathbb{Z}_{3}}\boxtimes\mathrm{Fib}$. Besides $\operatorname{Vect}_{\mathbb{Z}_{3}}\boxtimes\mathrm{Fib}$ and $\operatorname{Vect}$, which correspond to $\smash{\mathcal{P}(3)\otimes\textsl{VF}_{24}^{\natural}}$ and $\smash{V^{\natural}}$, respectively, there are the fusion subcategories $\operatorname{Vect}_{\mathbb{Z}_{3}}$ and $\mathrm{Fib}$. We call the corresponding vertex operator subalgebras predicted by 7.3 $U$ and $W$, respectively. There are thus the following Hasse diagrams: